# The Generation Phases of Flow Matching: a Denoising Perspective

### v1

@@ 1 Introduction
Flow matching (FM, Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu et al., 2023) and diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021) have achieved state-of-the-art results in generating images, videos, audio, and even text, where they are able to produce content that is virtually indistinguishable from human-produced ones. Research in the field remains extremely active, with many important questions still open: improving sample quality, making training and inference more efficient (Karras et al., 2022; Rombach et al., 2022) and, perhaps most importantly, understanding why current generative models perform so well.
Despite their striking successes, the precise mechanisms that make current generative models so effective still remain elusive. Identifying those is critical to improve them, and several leads have been proposed: analyzing the behaviour of the generative process across time (Biroli et al., 2024), exploiting connections with the exact minimizer of the training loss (Kamb and Ganguli, 2025; Niedoba et al., 2025), explaining memorisation and generalisation (Kadkhodaie et al., 2024; Sclocchi et al., 2025), or the impact of optimization procedures (Wu et al., 2025). Some of these studies, of theoretical nature, have provided elegant interpretations (e.g. the “target stochasticity” of Vastola, 2025, to explain generalisation), but were later questioned by empirical findings (Bertrand et al., 2025). This highlights the need for carefully designed empirical frameworks that can probe such theories and, in doing so, guide the development of better methods through deeper theoretical understanding.
@@
    We exploit the generality of the denoising framework, that allows building new models in a principled manner, with similar generation quality (FID-wise) but different generation behaviours. We highlight the importance of the intermediate stage for generation by learned models, a stage that is not revealed by the closed-form optimal velocity.
The structure of the paper is the following: Section 2 provides an introduction to flow matching; Section 3 lays down the theoretical equivalence between flow matching and denoising; Section 4 is dedicated to related works. Section 5 provides an empirical probe of the theoretical equivalence between denoising and generation. Section 6.1 leverages our framework to design relevant modifications of the generative process. Finally, Section 6 provides new insights on the phases of generation and their nature.
2 Background on flow matching
@@
$\min_{\theta}\,\mathbb{E}_{\begin{subarray}{c}t\sim\mathcal{U}[0,1],\\ x_{0}\sim p_{0},\\ x_{1}\sim p_{1}\end{subarray}}\left[\|v_{\theta}(x_{t},t)-(x_{1}-x_{0})\|^{2}\right],$ (2)
where $x_{t}:=(1-t)x_{0}+tx_{1}$ is the linear interpolation between $x_{0}$ and $x_{1}$. It is well-known that the solution $v^{\star}$ to this minimization problem (over all measurable functions) is given by a conditional expectation:
@@
In practice, sampling from $p_{1}$ in (2) is impossible, and points $x_{1}$ are instead drawn from a dataset $x^{(1)},\ldots,x^{(n)}$ of samples from $p_{1}$. Effectively, $p_{1}$ is replaced in (2) by the empirical measure $\hat{p}_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x^{(i)}}$. This has an important consequence: the minimizer of (2), when $p_{1}$ is replaced by $\hat{p}_{1}$, admits a closed-form $\hat{v}^{\star}$ (Gu et al., 2025; Bertrand et al., 2025):
@@
Paradoxically, using this closed-form velocity for generation can only reproduce samples from the training set. At the same time, this target velocity exhibits different behaviours across time: at $t=0$, it points towards the dataset mean, while as $t\to 1$ it points towards a single training point. Thus, understanding why trained FM models work comes down to determining, qualitatively and quantitatively, at which times $t$ and points $x$ the models must deviate from the closed-form in order to generate new samples, still consistent with the data distribution.
@@ 3.1 From flow matching to denoising
namely the Minimum Mean Square Error estimator of the clean image $x_{1}$ given the noisy observation $x_{t}$ at time $t$. Thus, any optimally trained FM model naturally yields an optimal denoiser. Building on (5), several works on flow-matching for image restoration define a denoiser at time $t$ as $D_{t}:x_{t}\mapsto x_{t}+(1-t)\,v(x_{t},t)$, see e.g. Zhang et al. (2024); Pokle et al. (2024); Martin et al. (2025).
@@ 3.2 From denoisers to velocities: the denoising toolkit
Generative denoisers differ from classical denoisers $D_{\sigma}$, parameterized by a noise level $\sigma$ and taking inputs of the form $x_{\sigma}=x+\sigma\varepsilon$. These two forms are equivalent up to rescaling. If the denoiser is parameterized by a noise level but the input is given as $x_{t}=tx+(1-t)\varepsilon$, one can define $x_{\sigma}=\frac{x_{t}}{t}$ and apply the classical denoiser $D_{\sigma}$ with $\sigma=\frac{1-t}{t}$ to approximately recover $x$. Conversely, if the input is of the form $x_{\sigma}=x+\sigma\varepsilon$, setting $x_{t}=\frac{x_{\sigma}}{1+\sigma}$ and $t=\frac{1}{1+\sigma}$, a generative denoiser $D_{t}$ can be applied to $x_{t}$ to recover $x$. The mapping from $\sigma$ to $t$ is a bijection between $0,\infty)$ and $(0,1]$.
@@ 3.3 Denoising losses $\mathcal{L}$
We present three types of denoising losses, all expressed for a generative denoiser $D:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$, taking as input $x_{t}=(1-t)x_{0}+tx_{1}$, with clean image $x_{1}\sim p_{1}$ and noise $x_{0}\sim\mathcal{N}(0,I_{d})$. Detailed derivations are provided in Appendix A.
@@
Compared to the FM loss $\mathcal{L}_{\text{FM}}$, classical denoising therefore differs in two important ways: (i) the loss includes a weight $w_{t}^{\text{classic}}:=\mathbf{1}_{[(1+\sigma_{\max})^{-1},1]}(t)\cdot t^{-2}$; (ii) the range of $t$ is truncated to $1/(1+\sigma_{\max}),1]$, which covers the full interval $[0,1]$ only if $\sigma_{\max}=\infty$. In other words, classical denoisers cannot handle small times in FM except if trained with unbounded noise levels. In practice, we set $\sigma_{\max}=19$ so that [^2] In traditional denoising, models are usually trained with noise level at most $\sigma=100/255\simeq 0.4$. $t_{\text{min}}=1/(1+\sigma_{\max})=0.05$.
@@
More general weightings in denoising losses. The three losses considered above emphasize opposite time intervals: $\mathcal{L}_{\mathrm{FM}}$ stresses large $t$ (lightly corrupted inputs), while $\mathcal{L}_{\text{classic}}$ stresses small $t$ (highly corrupted inputs). This is expressed by the time-dependent weightings $w_{t}$ in the loss $\mathbb{E}_{t,x_{0},x_{1}}\left[w_{t}\|D(x_{t},t)-x_{1}\|^{2}\right].$ As part of our methodology, in Section 6 we will also explore handcrafted weightings, allowing us to probe the generation process.
@@ 3.4 Equivalence in the optimal setting
A critical point is that all the above losses share the same minimizer in $L^{2}(\mathbb{R}^{d}\times[0,1])$, namely the MMSE denoiser $D^{\star}(x_{t},t)=\mathbb{E}[x_{1}|x_{t},t]$. However, in practice, minimization is restricted to a parametric function class $\mathcal{C}$ parametrized by a given neural architecture, optimization algorithms may be imperfect, and thus different loss weightings $w_{t}$ can lead to different numerical solutions.
@@ 3.5 Parametrizations of the denoisers and velocities
We now describe two parametrization classes $\mathcal{C}$ for the denoisers, which are used in the minimization of the losses $\mathcal{L}$ defined in Section 3.3. In all cases, $N^{\theta}$ denotes a neural network, with parameters $\theta$ belonging to some set $\Theta$.
@@
A key feature of this parametrization is that it enforces at $t=1$, $D_{t}={\rm Id}$, meaning the denoiser leaves its input unchanged which is the expected behavior (no noise in the input). It also matches the parametrization of standard flow matching models, where the velocity field is directly parametrized by a neural network. The different losses and parametrizations considered are summarized in Table 1.
4 Related works
Weighting strategies in denoising score matching.
We now discuss related works within the diffusion framework, which is known to be equivalent to flow matching when the source distribution $p_{0}$ is Gaussian (Gao et al., 2025). For simplicity, we consider the variance-preserving diffusion process where $x_{t}$ is obtained from a clean sample $x_{0}$ and a Gaussian noise $\varepsilon\sim\mathcal{N}(0,I_{d})$ as $x_{t}:=\alpha_{t}x_{0}+\sigma_{t}\varepsilon$ with $\alpha_{t}^{2}+\sigma_{t}^{2}=1$. Note that we use the standard diffusion notations in this section only, meaning that the sampling process evolves from $t=T$ (noise) to $t=0$ (clean image), as opposed to the flow matching convention where $t$ evolves from 0 to $1$. Regarding parametrization, most diffusion implementations train a network to predict the noise $\varepsilon$. Some papers (Salimans and Ho, 2022; Hang et al., 2023) also study $v$-prediction (i.e. $\mathcal{C}_{\mathrm{I+NN}}$) or $x$-prediction (i.e. $\mathcal{C}_{\mathrm{NN}}$).
Regarding weighting, most diffusion papers follow the formulation from Ho et al. (2020) which uses an unweighted loss in the $\varepsilon$-prediction, i.e. $\mathbb{E}_{t,x_{0},\varepsilon}[\|\varepsilon-\varepsilon^{\theta}(x_{t},t)\|^{2}]$, corresponding in the denoising framework to the weighting ${\alpha_{t}^{2}}/{\sigma_{t}^{2}}$, equal to the signal-to-noise ratio. A line of work is devoted to crafting loss weightings. Salimans and Ho (2022) propose the weighting $\max(\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}},1)$ while Yu et al. (2024) use weighting $\frac{\alpha_{t}}{\sigma_{t}}$. Both approaches assign greater importance to large noise levels, arguing that these correspond to more difficult denoising tasks and play a crucial role in error propagation during sampling. Interpreting diffusion training as multitask optimization, Hang et al. (2023) observe conflict between different timesteps objectives and suggest weighting $\min(\frac{\alpha_{t}^{2}}{\sigma_{t}^{2}},\gamma)$ to avoid putting too much weight on the low noise levels, identified as an easy denoising task. In contrast, Choi et al. (2022) introduce the P2 weighting that puts more weight on the intermediate times, hypothesizing that perceptual features emerge during this content phase, as opposed to the early coarse phase or the late cleanup phase, where little noise remains in the image.
Rather than hypothesizing which timesteps are important or directly relating the weighting strategy to FID performance, we introduce an intermediate test metric–the PSNR evaluated at each timestep–which enables a more fine-grained analysis of how different timesteps contribute to generation quality. Similar to Choi et al. (2022), we find that effective denoising at intermediate times is crucial for high-quality generation.
Unifying perspectives on loss weightings.
Kingma and Gao (2023) provide a unifying view on a wide range of loss weightings (including FM and the previous mentioned ones) by rewriting them as differently weighted ELBO formulations. Kumar et al. (2025) also lay down the different losses (noise prediction, score prediction, etc) and the loss weighting they induce.
Linking diffusion and denoising.
Beyond standard diffusion, Delbracio and Milanfar (2023) study a broader “degradation-to-clean” setting (which is not limited to Gaussian denoising) and pick a standard denoising loss (i.e. weighting $w_{t}^{\mathrm{den}}=1$, parametrization $\mathcal{C}_{\mathrm{NN}}$). For diffusion, they report good results on CelebA ($64\times 64$), yet below state-of-the-art levels, which is consistent with our observations. Leclaire et al. (2025) also provide a synthetic overview of the connections between diffusion models and classical additive Gaussian denoising, interpreting diffusion as an iterative “Noising-Relaxed Denoising” process to better understand the role of noise schedules.
Analysis of the generation phases.
One line of research focuses directly on the closed-form, either theoretically (Biroli et al., 2024) or empirically (Bertrand et al., 2025). Both works highlight a critical time, at early timesteps, beyond which the closed-form points towards a single training example, while the trained models begin to deviate from it.
Another group of works seeks to understand the reasons for generalisation by characterizing the velocities or scores that are actually learned. Kadkhodaie et al. (2024); Niedoba et al. (2025); Kamb and Ganguli (2025) analyze the effective receptive field of trained models and show that it evolves from global (at high noise levels) to local (at low noise levels). Niedoba et al. (2025); Kamb and Ganguli (2025) both argue that what is actually learned by the models is a patch-wise version of the closed-form, with patch size determined by the model’s receptive field: they demonstrate that this new formulation can predict in certain cases, without training, the samples learned by the model.
Beyond works focusing on the closed-form and its approximation, Sclocchi et al. (2025) connect the diffusion generative process to the learning of high- and low-level features: they specifically show that there exists a critical time at which image class is determined, while low-level features evolve smoothly throughout the generation process.
Finally, some works explain generalisation by imperfect or early-stopped optimization (Wu et al., 2025; Bonnaire et al., 2025; Favero et al., 2025).
@@ 5.1 Denoising and generative metrics
As a first investigation, we evaluate models trained using the different couples of denoising losses/parametrizations ($\mathcal{L}$, $\mathcal{C}$) on CIFAR-10 ($\times32$, Krizhevsky and Hinton, 2009) and CelebA-64 ($\times64$, Yang et al., 2015). All models share the same architecture and training hyperparameters, borrowed from the standard FM training (full details in Appendix B). We also train 10 independent denoisers with $\mathcal{L}_{\mathrm{den}}$, each trained only on the time interval $[i/10,(i+1)/10]$ for $i\in\llbracket 0,9\rrbracket$, yielding a velocity field defined in a piecewise manner over time; the resulting model is denoted “10-denoisers”.
Figure 2 displays the performance of our trained models both in denoising and in generation. Denoising at noise level/time $t$ is evaluated using the Peak Signal-to-Noise Ratio (PSNR) between denoiser output $D_{t}(x_{t})$ and clean sample $x_{1}\sim p_{1}$, averaged on $1000$ test images. Generation is evaluated using the Fréchet Inception Distance (FID, Heusel et al., 2017) between 10k test samples and 10k generated samples. First, Figure 2 confirms that, although all denoising losses $\mathcal{L}$ are equivalent in terms of the optimal denoiser they promote, both the choices of $\mathcal{L}$ and $\mathcal{C}$ impact performance in practice [^3] We provide in Appendix C a table with additional tested weightings, showing the same trends.. For every loss except the $\mathcal{L}_{\text{classic}}$ (for which the generated images are of such poor quality that the corresponding FID values are not meaningful), the residual parametrization $\mathcal{C}_{\mathrm{I+NN}}$ consistently outperforms the plain parametrization $\mathcal{C}_{\text{NN}}$, supporting the hypothesis that explicitly enforcing $D_{t}={\rm Id}$ introduces a beneficial implicit bias. Second, the models with lower FID also have better PSNRs at all noise levels, except for 10-denoisers which has worse PSNR than standard flow matching for $t<0.9$. The best performance is obtained with the FM loss under the parametrization $\mathcal{C}_{\mathrm{I+NN}}$ (i.e. the standard flow matching approach). On the contrary, the models trained with the classical loss ($w_{t}^{\text{classic}}=t^{-2}$) and the unweighted denoising loss ($w_{t}^{\text{den}}=1$), both motivated only by denoising considerations, obtain not only poorer generation performance, but also poorer denoising performance. Combined with the good performance of the 10-denoisers, this suggests that training a single network (taking $t$ as a parameter) to denoise at every noise level, if not counterbalanced with an appropriate use of weights, is detrimental to performance. Perhaps surprisingly, the FM weights $w_{t}^{\text{FM}}=1/(1-t)^{2}$ turn out to be the most efficient for denoising, although they put more emphasis on accurate denoising at low noise level ($t$ close to $1$), which one may think of as an easy task.

### v2

@@ 1 Introduction
Flow matching (FM, lipman_flow_2022; albergo2023stochasticinterpolant; liu2023rectifiedflow) and diffusion models (sohlDickstein2015diffusion; Ho2020; Song2021) have achieved state-of-the-art results in generating images, videos, audio, and even text, where they are able to produce content that is virtually indistinguishable from human-produced ones. Research in the field remains extremely active, with many important questions still open: improving sample quality, making training and inference more efficient (karras2022elucidating; rombach2022high) and, perhaps most importantly, understanding why current generative models perform so well.
Despite their striking successes, the precise mechanisms that make current generative models so effective still remain elusive. Identifying those is critical to improve them, and several leads have been proposed: analyzing the behaviour of the generative process across time (biroli2024dynamical), exploiting connections with the exact minimizer of the training loss (kamb2024analytic; niedoba2024towards), explaining memorisation and generalisation (kadkhodaie2024generalisation; sclocchi2025phase), or the impact of optimization procedures (wu2025taking). Some of these studies, of theoretical nature, have provided elegant interpretations (e.g. the “target stochasticity” of vastola2025generalisation, to explain generalisation), but were later questioned by empirical findings (bertrand2025closed). This highlights the need for carefully designed empirical frameworks that can probe such theories and, in doing so, guide the development of better methods through deeper theoretical understanding.
@@
    We exploit the generality of the denoising framework, that allows building new models in a principled manner, with similar generation performance but different generation behaviours. We highlight the importance of the intermediate stage for generation by learned models, a stage that is not revealed by the closed-form optimal velocity.
The structure of the paper is the following: Section 2 provides an introduction to flow matching and related works; Section 3 lays down the theoretical equivalence between flow matching and denoising, which Section 4 empirically attests. Section 5 leverages our framework to design relevant modifications of the generative process. Finally, Section 6 provides new insights on the phases of generation and their nature.
2 Preliminaries
Background on flow matching
@@ 2 Preliminaries
$\min_{\theta}\,\mathbb{E}_{t\sim\mathcal{U}[0,1],\,x_{0}\sim p_{0},\,x_{1}\sim p_{1}}\left[\|v_{\theta}(x_{t},t)-(x_{1}-x_{0})\|^{2}\right],$ (2)
where $x_{t}:=(1-t)x_{0}+tx_{1}$ is the linear interpolation between $x_{0}$ and $x_{1}$. It is well-known that the solution $v^{\star}$ to this problem (over all measurable functions) is given by a conditional expectation:
@@
In practice, sampling from $p_{1}$ in (2) is impossible, and points $x_{1}$ are instead drawn from a dataset $x^{(1)},\ldots,x^{(n)}$ of samples from $p_{1}$. Effectively, $p_{1}$ is replaced in (2) by the empirical measure $\hat{p}_{1}=\frac{1}{n}\sum_{i=1}^{n}\delta_{x^{(i)}}$. This has an important consequence: the minimizer of (2), when $p_{1}$ is replaced by $\hat{p}_{1}$, admits a closed-form $\hat{v}^{\star}$ (gu2025memorisation; bertrand2025closed):
@@
Paradoxically, using this closed-form velocity for generation can only reproduce samples from the training set. At the same time, this target velocity exhibits different behaviours across time: at $t=0$, it points towards the dataset mean, while as $t\to 1$ it points towards a single training point. Thus, understanding why trained FM models work comes down to determine, qualitatively and quantitatively, at which times $t$ and points $x$ the models must deviate from the closed-form in order to generate new samples, still consistent with the data distribution.
Related works
First, flow matching with Gaussian $p_{0}$ is known to be equivalent to diffusion, for which there is an abundant literature about weighting the loss function across time (hang2023efficient; kingma2023understanding); in diffusion, kumar2025loss lays down the different losses (noise prediction, score prediction, etc) and the loss weighting they induce. leclaire2025backward interprets diffusion as iterative “Noising-Relaxed Denoising” to better understand the noise schedules. Several works highlight the existence of distinct phases in the generative process (biroli2024dynamical; bertrand2025closed), which sclocchi2025phase connect to learning of high- and low-level features; a generalisation followed by a memorisation phase also seems to appear in training (bonnaire2025diffusion). A line of work seeks to understand the reasons for generalisation in generative models, characterizing the velocities/scores that are actually learned (kadkhodaie2024generalisation; kamb2024analytic; niedoba2024towards). We refer the reader to Appendix A for a detailed discussion of related works.
@@ 3.1 From flow matching to denoising
namely the Minimum Mean Square Error estimator of the clean image $x_{1}$ given the noisy observation $x_{t}$ at time $t$. Thus, any optimally trained FM model naturally yields an optimal denoiser. Building on (5), several works on flow-matching for image restoration define a denoiser at time $t$ as $D_{t}:x_{t}\mapsto x_{t}+(1-t)\,v(x_{t},t)$, see e.g. zhang2024flow; pokle2023training; martin2024pnp.
@@ 3.2 From denoisers to velocities: the denoising toolkit
Generative denoisers differ from classical denoisers $D_{\sigma}$, parameterized by a noise level $\sigma$ and taking inputs of the form $x_{\sigma}=x+\sigma\varepsilon$. These two forms are equivalent up to rescaling. If the denoiser is parameterized by a noise level but the input is given as $x_{t}=tx+(1-t)\varepsilon$, one can define $x_{\sigma}=\frac{x_{t}}{t}$ and apply the classical denoiser $D_{\sigma}$ with $\sigma=\frac{1-t}{t}$ to approximately recover $x$. Conversely, if the input is of the form $x_{\sigma}=x+\sigma\varepsilon$, setting $\tilde{x}=\frac{x_{\sigma}}{1+\sigma}$ and $t=\frac{1}{1+\sigma}$, a generative denoiser $D_{t}$ can be applied to $\tilde{x}$ to recover $x$. The mapping from $\sigma$ to $t$ is a bijection between $0,\infty)$ and $(0,1]$.
@@ 3.3 Denoising losses $\mathcal{L}$
We present three types of denoising losses, all expressed for a generative denoiser $D:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$, taking as input $x_{t}=(1-t)x_{0}+tx_{1}$, with clean image $x_{1}\sim p_{1}$ and noise $x_{0}\sim\mathcal{N}(0,I_{d})$.
@@
Compared to the FM loss $\mathcal{L}_{\text{FM}}$, classical denoising therefore differs in two important ways: (i) the loss includes a weight $w_{t}^{\text{classic}}:=\mathbf{1}_{[(1+\sigma_{\max})^{-1},1]}(t)\cdot t^{-2}$; (ii) the range of $t$ is truncated to $1/(1+\sigma_{\max}),1]$, which covers the full interval $[0,1]$ only if $\sigma_{\max}=\infty$. In other words, classical denoisers cannot handle very low SNR regimes except if trained with unbounded noise levels. In practice, we set $\sigma_{\max}=19$ so that [^2] In traditional denoising, models are usually trained with noise level at most $\sigma=100/255\simeq 0.4$. $t_{\text{min}}=1/(1+\sigma_{\max})=0.05$.
@@
More general weightings in denoising losses. The three losses considered above emphasize opposite time intervals: $\mathcal{L}_{\mathrm{FM}}$ stresses large $t$ (lightly corrupted inputs), while $\mathcal{L}_{\text{classic}}$ stresses small $t$ (highly corrupted inputs). This is expressed by the time-dependent weightings $w_{t}$ in the loss $\mathbb{E}_{t,x_{0},x_{1}}\left[w_{t}\|D(x_{t},t)-x_{1}\|^{2}\right].$ As part of our methodology, in Section 6 we will also explore handcrafted weightings, allowing us to probe the generation process. kingma2023understanding provide similar derivations and additional weightings.
@@ 3.4 Equivalence in the optimal setting
A critical point is that all the above losses share the same minimizer in $L^{2}(\mathbb{R}^{d}\times[0,1])$, namely the MMSE denoiser $D^{\star}(x_{t},t)=\mathbb{E}[x_{1}|x_{t},t]$. However, in practice, minimization is restricted to a parametric function class $\mathcal{C}$, optimization algorithms may be imperfect, and thus different loss weightings can lead to different numerical solutions.
@@ 3.5 Parametrizations of the denoisers and velocities
We now describe two parametrization classes $\mathcal{C}$ for the denoisers, which are used in the minimization of the losses $\mathcal{L}$ defined in the previous subsection. In all cases, $N^{\theta}$ denotes a neural network, with parameters $\theta$ belonging to some set $\Theta$.
@@
A key feature of this parametrization is that it enforces at $t=1$, $D_{t}={\rm Id}$, meaning the denoiser leaves its input unchanged – which is the expected behavior (no noise in the input). It also matches the parametrization of standard flow matching models, where the velocity field is directly parametrized by a neural network. We can already hypothesize that this implicit bias may benefit the model. The different losses and parametrizations considered are summarized in Table 1.
(b) FID on 50k train images (lower is better).
@@ 4.1 Denoising and generative metrics
As a first investigation, we evaluate models trained using the different couples of denoising losses/parametrizations ($\mathcal{L}$, $\mathcal{C}$) on CIFAR-10 ($\times32$, krizhevsky2009learning) and CelebA-64 ($\times64$, yang2015facial). All models share the same architecture and training hyperparameters, borrowed from the standard FM training (full details in Appendix C). We also train 10 independent denoisers with $\mathcal{L}_{\mathrm{den}}$, each trained only on the time interval $[i/10,(i+1)/10]$ for $i\in\llbracket 0,9\rrbracket$, yielding a velocity field defined in a piecewise manner over time; the resulting model is denoted “10-denoisers”.
Figure 2 displays the performance of our trained models both in denoising and in generation. Denoising at noise level/time $t$ is evaluated using the Peak Signal-to-Noise Ratio (PSNR) between denoiser output $D_{t}(x_{t})$ and clean sample $x_{1}\sim p_{1}$, averaged on $1000$ test images. Generation is evaluated using the Fréchet Inception Distance (FID, heusel2017) between 50k train samples and 50k generated samples. First, Figure 2 confirms that, although all denoising losses $\mathcal{L}$ are equivalent in terms of the optimal denoiser they promote, both the choices of $\mathcal{L}$ and $\mathcal{C}$ impact performance in practice [^3] We provide in Appendix D a table with additional tested weightings, showing the same trends.. For every loss except the $\mathcal{L}_{\text{classic}}$ (for which the generated images are of such poor quality that the corresponding FID values are not meaningful), the residual parametrization $\mathcal{C}_{\mathrm{I+NN}}$ consistently outperforms the plain parametrization $\mathcal{C}_{\text{NN}}$, supporting the hypothesis that explicitly enforcing $D_{t}={\rm Id}$ introduces a beneficial implicit bias. Second, the models with lower FID also have better PSNRs at all noise levels, except for 10-denoisers which has worse PSNR than standard flow matching for $t<0.9$. The best performance is obtained with the FM loss under the parametrization $\mathcal{C}_{\mathrm{I+NN}}$ (i.e. the standard flow matching approach). On the contrary, the models trained with the classical loss ($w_{t}^{\text{classic}}=t^{-2}$) and the unweighted denoising loss ($w_{t}^{\text{den}}=1$), both motivated only by denoising considerations, obtain not only poorer generation performance, but also poorer denoising performance. Combined with the good performance of the 10-denoisers, this suggests that training a single network (taking $t$ as a parameter) to denoise at every noise level, if not counterbalanced with an appropriate use of weights, is detrimental to performance. Perhaps surprisingly, the FM weights $w_{t}^{\text{FM}}=1/(1-t)^{2}$ turn out to be the most efficient for denoising, although they put more emphasis on accurate denoising at low noise level ($t$ close to $1$), which one may think of as an easy task.